### v1

@@ 1 Introduction
The combined complexity of workload diversity and hardware heterogeneity complicates the development of a comprehensive attention solution. Currently, each system implements a specialized attention solution based on a subset of these characteristics, leading to high maintenance overhead and potential inefficiencies. To address these challenges, we introduce FlashInfer  [^1] Project page: https://flashinfer.ai/, a code-generation based attention engine designed to accelerate attention computation in LLMs. Our approach incorporates several key designs:
@@ 3.2.3 JIT Compiler for Attention Variants
Figure 5 shows how FlashInfer maps FlashSigmoid’s  Ramapuram et al. (2024) attention specification to different parts of FlashInfer’s CUDA templates. Attention specificiation accepts a piece of CUDA code to define the variant functors, such design also enables user to use advanced PTX instructions  [^2] https://docs.nvidia.com/cuda/parallel-thread-execution/ or even their own libraries. The JIT compiler generates the CUDA code by inserting the variant class and other information into the template, and the generated CUDA code is compiled with PyTorch’s JIT compiler  [^3] https://pytorch.org/tutorials/advanced/cpp_extension.html#jit-compiling-extensions and registered as a custom operator  [^4] https://pytorch.org/tutorials/advanced/custom_ops_landing_page.html.
@@ 3.3.1 Load-balanced Scheduling
FlashInfer guarantees both attention and contraction kernels are compatible with CUDAGraphs  Gray (2019); Nguyen et al. (2021). Both attention and contraction kernels use persistent kernel and the grid size is fixed once compiled, which means the kernel is launched with the same grid size for each generation step. We set the fixed offset for each section of the workspace buffer to store partial outputs and plan information to make sure the pointers passed to the kernel are the same for each generation step, meeting the requirement of CUDAGraphs (see Appendix D.1 for details).
@@ 3.4 Programming Interface
The plan function activates the dynamic scheduler by processing sequence length data to generate load-balanced scheduling plans. These plans are cacheable, allowing reuse across operators with matching sequence length specs, such as all decode attentions in a generation step. The run function executes the attention computation using inputs of query, key, value, and cached plan data, outputting the attention results. CUDAGraph can capture calls to run functions and compile the entire attention generation step into a single graph. However, plan function is not captured by CUDAGraph because it’s on CPU.
@@ 4.1 End-to-end LLM serving performance
We evaluate FlashInfer with SGLang v0.3.4 Zheng et al. (2023b) and compare its performance against two settings: SGLang with Triton v3.0 Tillet et al. (2019) and TensorRT-LLM v0.13.0 NVIDIA (2023a). The latter is a leading LLM serving engine optimized for NVIDIA GPUs; however, its attention kernels are closed-source, which limits transparency and potential for community-driven improvements. To ensure a comprehensive evaluation, we employ two datasets: the widely-used ShareGPT dataset  [^5] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered/resolve/main/ShareGPT_V3_unfiltered_cleaned_split.json and a synthetic workload (Variable) with sequence lengths uniformly distributed between 512 and 2048 tokens. We measure the TTFT(time-to-first-token) and ITL(inter-token-latency) under latency-sensitive online serving settings, the request rate is adjusted to maintain P99 TTFT below 200ms.
Figure 7: Medium Inter-Token-Latency (ITL) and medium Time-To-First-Token (TTFT) of SGLang integrated with FlashInfer and Triton, compared to TensorRT-LLM.
Figure 7 shows the ITL and TTFT measured on both Llama 3.1  Dubey et al. (2024) 8B (on 1xH100) and Llama 3.1 70B (on 4xH100) models. Compared to SGLang with Triton backend, FlashInfer backend shows consistent speedup in all settings. Compared to TRTLLM, FlashInfer demonstrate weaker performance on ShareGPT (especially on TTFT for Llama 3.1 70B), but the performance of SGLang + FlashInfer is similar to TRTLLM on Variable workload. The performance gap between SGLang and TRTLLM is not only due to the attention kernels, but also the difference in other kernels (e.g. allreduce) and system design. Overall, FlashInfer backend shows competitive performance on LLM serving benchmarks.
@@ 6 Conclusion and Future Work
In this paper, we present FlashInfer, an versatile and efficient attention engine for LLM serving. We propose a unified block-sparse storage and composable formats for memory efficiency, JIT compilation for customization and load-balanced scheduler for input dynamism. We evaluate FlashInfer’s performance across diverse inference scenarios, showing strong performance in kernel-level and end-to-end LLM serving metrics. In the future, we plan to explore compiling higher-level DSLs  Wu et al. (2024); He et al. (2024) to attention specifications in FlashInfer, as well as code generation to other backends  Ozen (2024); Spector et al. (2024); Tillet et al. (2019). The FlashInfer project is open-source and available at https://github.com/flashinfer-ai/flashinfer.
@@
We would like to thank members of LMSYS ORG, UW Syslab and SAMPL research group, CMU Catalyst group and for their useful feedback and discussions. We thank Yaxing Cai, Junru Shao, Lianmin Zheng, Ying Sheng, Liangsheng Yin, Lily Liu, Woosuk Kwon, Cody Yu, Ray Wan, Bowen Wang, Pavani Majety, Elfie Guo, Travis Addair, Cuimi Guo for their help in integrating FlashInfer into LLM serving frameworks. We also thank Zhuoming Chen, Lesheng Jin, Antoni Baum, Kaichao You, Simon Mo, Ke Bao, Byron Hsu, Zhiqiang Xie, Haofeng Huang, Sirui Lu, Henry Xiao, Chi-Chih Chang, Yilong Zhao, Size Zheng, Bohan Hou, Yang Yu, Nandor Licker, Tsu Bin, Hieu Pham, Horace He, Vijay Thakkar, Yuxian Qiu, Freddy Qi, June Yang, Bing Xu, Anxhelo Xhebraj, Evghenii Gaburov, Bastian Hagedorn and all community contributors for their input and feedback on the FlashInfer project. Researchers from CMU are supported by the National Science Foundation (NSF) under award CNS-2211882. Researchers from UW are supported in part by NSF under award CCF-1518703, and by ACE and PRISM, two of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA; Zihao Ye is supported by NVIDIA Graduate Fellowship. Luis Ceze is supported by the Lazowska Endowed Professorship. The opinions and conclusions in this paper do not reflect the views of these funding agencies.
@@ D.2 Split-K Writethrough Optimizations
CUTLASS  (Thakkar et al., 2023) implements deterministic turnstile accumulation (for GEMM), which eliminates the need for a workspace buffer for partial outputs by using semaphores to ensure a deterministic reduction order, the same idea also applies to attention composition. We leave this implementation as future work.

### v2

@@ 1 Introduction
The combined complexity of workload diversity and hardware heterogeneity complicates the development of a comprehensive attention solution. Currently, each system implements a specialized attention solution based on a subset of these characteristics, leading to high maintenance overhead and potential inefficiencies. To address these challenges, we introduce FlashInfer, a code-generation based attention engine designed to accelerate attention computation in LLMs. Our approach incorporates several key designs:
@@ 3.2.3 JIT Compiler for Attention Variants
Figure 5 shows how FlashInfer maps FlashSigmoid’s  Ramapuram et al. (2024) attention specification to different parts of FlashInfer’s CUDA templates. Attention specificiation accepts a piece of CUDA code to define the variant functors, such design also enables user to use advanced PTX instructions  [^1] https://docs.nvidia.com/cuda/parallel-thread-execution/ or even their own libraries. The JIT compiler generates the CUDA code by inserting the variant class and other information into the template, and the generated CUDA code is compiled with PyTorch’s JIT compiler  [^2] https://pytorch.org/tutorials/advanced/cpp_extension.html#jit-compiling-extensions and registered as a custom operator  [^3] https://pytorch.org/tutorials/advanced/custom_ops_landing_page.html. We also support compiling to other runtime systems through a framework agnostic DLPack  [^4] https://github.com/dmlc/dlpack interface.
@@ 3.3.1 Load-balanced Scheduling
FlashInfer guarantees both attention and contraction stage are compatible with CUDAGraphs  Gray (2019); Nguyen et al. (2021). Both attention and contraction stage use persistent kernel and the grid size is fixed once compiled, which means the kernel is launched with the same grid size for each generation step. We set the fixed offset for each section of the workspace buffer to store partial outputs and plan information to make sure the pointers passed to the kernel are the same for each generation step, meeting the requirement of CUDAGraphs (see Appendix D.1 for details). We merge the two stages into one persistent kernel, eliminating intra-kernel overhead.
@@ 3.4 Programming Interface
The plan function activates the dynamic scheduler by processing sequence length data to generate load-balanced scheduling plans. These plans are cacheable, allowing reuse across operators with matching sequence length specs, such as all decode attentions in a generation step. The run function executes the attention computation using inputs of query, key, value, and cached plan data, outputting the attention results. CUDAGraph can capture calls to run functions and compile the entire attention generation step into a single graph. However, plan function is not captured by CUDAGraph because it’s on CPU. The plan and run division is inspired by the Inspector-Executor (IE) model Mirchandaney et al. (1988); Saltz & Mirchandaney (1991); Ponnusamy et al. (1993), which is widely used for parallelizing irregular workloads.
@@ 4.1 End-to-end LLM serving performance
We evaluate FlashInfer with SGLang v0.3.4 Zheng et al. (2023b) and compare its performance against two settings: SGLang with Triton v3.0 Tillet et al. (2019). The latter is a leading LLM serving engine optimized for NVIDIA GPUs; however, its attention kernels are closed-source, which limits transparency and potential for community-driven improvements. To ensure a comprehensive evaluation, we employ two datasets: the widely-used ShareGPT dataset  [^5] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered/resolve/main/ShareGPT_V3_unfiltered_cleaned_split.json and a synthetic workload (Variable) with sequence lengths uniformly distributed between 512 and 2048 tokens. We measure the TTFT(time-to-first-token) and ITL(inter-token-latency) under latency-sensitive online serving settings, the request rate is adjusted to maintain P99 TTFT below 200ms.
Figure 7: Medium Inter-Token-Latency (ITL) and medium Time-To-First-Token (TTFT) of SGLang integrated with FlashInfer and Triton.
Figure 7 shows the ITL and TTFT measured on both Llama 3.1  Dubey et al. (2024) 8B (on 1xH100) and Llama 3.1 70B (on 4xH100) models. Compared to SGLang with Triton backend, FlashInfer backend shows consistent speedup in all settings.
@@ 5.4 LLM Serving Systems
6 Discussions
Currently, FlashInfer only supports the forward pass for attention computation. To extend FlashInfer and apply it to training would require developing customizable backward attention kernel templates, which we plan to explore in future work. Regarding the generality of FlashInfer, our approach decouples computation from tile scheduling, allowing for diverse tiling strategies such as FlashDecoding Hong et al. (2024) and Lean Attention Sanovar et al. (2024) by delegating scheduling to a runtime-scheduler rather than embedding it within the attention template. This design generalizes scheduling algorithms, as detailed in Algorithm 1, for load-balancing and wave-quantization reduction while targeting optimal GPU performance across architectures (e.g., FlashAttention2 Dao (2023) for Turing/Ampere/Ada and FlashAttention3 Shah et al. (2024) for Hopper). Although frameworks like Triton Tillet et al. (2019) offer GPU-agnostic interfaces, they often lag in adopting new hardware features; our template design space, expressed as $f_{\textrm{epilogue}}(\textrm{scan}(f_{\textrm{logits}}(f_{q}(Q)\cdot f_{k}(K)))\cdot f_{v}(V))$, covers most attention functions, including recent variants such as Multi-head Latent Attention (MLA) DeepSeek-AI et al. (2024) and the intra-attention component of Linear Attention Yang et al. (2024).
@@ 7 Conclusion and Future Work
In this paper, we present FlashInfer, an versatile and efficient attention engine for LLM serving. We propose a unified block-sparse storage and composable formats for memory efficiency, JIT compilation for customization and load-balanced scheduler for input dynamism. We evaluate FlashInfer’s performance across diverse inference scenarios, showing strong performance in kernel-level and end-to-end LLM serving metrics. In the future, we plan to explore compiling higher-level DSLs  Wu et al. (2024); He et al. (2024) to attention specifications in FlashInfer, as well as code generation to other backends  Ozen (2024); Spector et al. (2024); Tillet et al. (2019). The FlashInfer project is open source and available at https://github.com/flashinfer-ai/flashinfer, and has been deployed at scale in production-level systems.
@@
We thank anonymous MLSys reviewers for providing constructive comments, LMSYS ORG, UW Syslab and SAMPL research group, CMU Catalyst group for their useful feedback and discussions, Yaxing Cai, Junru Shao, Lianmin Zheng, Ying Sheng, Liangsheng Yin, Lily Liu, Woosuk Kwon, Cody Yu, Ray Wan, Bowen Wang, Pavani Majety, Elfie Guo, Travis Addair, Cuimi Guo for their help in integrating FlashInfer into LLM serving frameworks, Zhuoming Chen, Lesheng Jin, Antoni Baum, Kaichao You, Simon Mo, Ke Bao, Byron Hsu, Zhiqiang Xie, Haofeng Huang, Sirui Lu, Henry Xiao, Chi-Chih Chang, Yilong Zhao, Size Zheng, Bohan Hou, Yang Yu, Nandor Licker, Tsu Bin, Hieu Pham, Horace He, Vijay Thakkar, Yuxian Qiu, Freddy Qi, June Yang, Bing Xu, Anxhelo Xhebraj, Evghenii Gaburov, Bastian Hagedorn and all community contributors for their input and feedback on the FlashInfer project. This work was supported in part by NSF award CNS-2211882, and gifts from OctoAI, Qualcomm, and CMU opensource software fellowships. Researchers from UW are supported in part by NSF under award CCF-1518703, and by ACE and PRISM, two of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA; Zihao Ye is supported by the NVIDIA Graduate Fellowship. Luis Ceze is supported by the Lazowska Endowed Professorship. The opinions and conclusions in this paper do not reflect the views of these funding agencies.
@@ Appendix F FP8–FP16 Mixed-Precision Attention
Appendix G Additional Evaluation
In this section, we present additional evaluation results to further validate the performance, scalability, and robustness of FlashInfer across diverse experimental conditions.
G.1 Comparison with FlexAttention
We compare FlashInfer and FlexAttention He et al. (2024) on different attention variants using the AttentionGym PyTorch-Labs (2024) benchmark on NVIDIA H100 80GB SXM. We evaluated with batch size $16$, number of heads $16$ and head dim $128$, the CUDA version and the Triton version were fixed to 12.4 and 3.2, respectivelyrespectively. Tables 1 to 4 show the performance of FlashInfer and FlexAttention in TFLOPS/s, where higher numbers mean better performance. Across all four scenarios and a range of sequence lengths, FlashInfer consistently outperforms FlexAttention, with especially large gains at longer sequence lengths. The better performance is mainly due to the usage of Hopper microarchitecture’s advanced features (such as warp specialization and TMA), and CUTLASS’s fine-grained resource control (at register-level rather than tile-level) over Triton. Note that these gaps will be alleviated once Triton fully supports these features.
Table 1: Causal Attention
Seq Length FlexAttention FlashInfer
512 209.11 250.454
1024 294.53 406.554
2048 376.90 487.236
4096 421.00 548.388
8192 441.26 587.903
16384 453.57 612.259
Table 2: Attention with Logits SoftCap
Seq Length FlexAttention FlashInfer
512 241.51 336.487
1024 327.50 409.534
2048 379.57 468.769
4096 403.39 489.667
8192 407.82 515.573
16384 409.89 520.935
Table 3: ALiBi Bias Press et al. (2022)
Seq Length FlexAttention FlashInfer
512 253.22 403.899
1024 344.70 500.220
2048 406.14 535.498
4096 426.13 561.324
8192 436.35 573.493
16384 434.86 578.005
Table 4: Sliding Window (window size = $1024$)
Seq Length FlexAttention FlashInfer
512 206.51 236.363
1024 292.25 374.108
2048 350.91 381.464
4096 368.45 384.998
8192 373.25 384.514
16384 367.91 380.506
G.2 Evaluation of Shared-Prefix Attention Kernels
We measure shared-prefix attention kernels with suffix length $128$. Table 5 shows the kernel latency under different shared prefix lengths, scenarios and batch sizes, where numbers are in microseconds (us), and “composable” means composable format while “single” means single format. The composable format benefits long prefixes (e.g., 32k) and large batch sizes (e.g., $64$). However, these speedups do not always yield proportional end-to-end gains because real-world shared prefix sizes tend to be smaller.
Table 5: Latency of Shared-Prefix Attention Kernels
Shared Prefix Length Composable (BS=16) Single (BS=16) Composable (BS=64) Single (BS=64)
1024 45.17 46.52 87.86 130.49
8192 88.67 226.57 125.76 931.75
32768 217.42 945.67 254.54 4090
G.3 Ablation Study on Variable Sequence Length and load-balancing scheduler
We conduct ablations on the effect of load-balancing scheduler (Section 3.3.1). Table 6 and 7 show the results for Llama 3.1-8B-Instruct running on an NVIDIA H100 SXM5 GPU with SGLang Zheng et al. (2023b) + FlashInfer (with and without load-balancing scheduler). We evaluate the inter-token latency (ITL, ms) and time-to-first-token (TTFT, ms) with three datasets: ShareGPT, variable sequence length with input lengths sampled from $U(512,2048)$ and output fixed at $256$, and variable sequence length with input lengths sampled from $U(4096,16384)$ and output fixed at $256$. “RR” in the tables means request rate.
Table 6: Load-balancing Scheduler Ablation Study (ITL)
Scenario w/ Load- Balancing w/o Load- Balancing Triton
ShareGPT (RR=16) 8.96 9.16 9.36
$U(512,2048)$ (RR=8) 8.21 8.42 8.49
$U(4096,16384)$ (RR=1) 8.63 13.89 11.08
Table 7: Load-balancing Scheduler Ablation Study (TTFT)
Scenario w/ Load- Balancing w/o Load- Balancing Triton
ShareGPT (RR=16) 39.05 39.42 52.92
$U(512,2048)$ (RR=8) 66.78 67.38 68.48
$U(4096,16384)$ (RR=1) 411.02 421.60 566.30
G.4 vLLM Integration Evaluation
We compare the vLLM with FlashInfer backend and its default backend with a fixed request rate of $16$, reporting throughput (tokens/s), inter-token latency (ITL, ms), and time-to-first-token (TTFT, ms) in Table 8. FlashInfer reduces ITL by aroudn 13% using fp8 KV-cache, but heavy Python overhead in vLLM integration (e.g. array operations) at host side causes minor regressions with bf16. Our future optimizations will address these in C++ and move the scheduler to device.
Table 8: vLLM Integration Evaluation
Backend Throughput Median ITL Median TTFT
Default (bf16) 6062.89 10.42 35.85
FlashInfer (bf16) 6065.41 10.63 36.60
Default (e4m3) 6015.86 12.56 39.74
FlashInfer (e4m3) 6020.32 10.92 37.93
G.5 Fine-Grained Block-Sparsity Evaluation
FlashInfer supports fine-grained block-sparse matrices, which is useful in many KV-Cache pruning algorithms. We measure the kernel performance on Quest Tang et al. (2024), a state-of-the-art long-context modeling algorithm, which uses fine-grained sparsity in KV-Cache. We compared the batch decoding attention kernel in Quest using FlashInfer and compared its performance to PyTorch SDPA and FlexAttention on an NVIDIA H100 SXM5 GPU with the configuration (block size 16, num_qo_heads $32$, num_kv_heads $32$, head_dim $128$). All latency values reported are in microseconds (us).
As shown in Table 9 to 11, FlashInfer demonstrates a considerable performance advantage, achieving up to a 20x speedup for long sequence lengths. Currently FlexAttention relies on large block size templates, while FlashInfer employs a sparse-row gathering strategy to leverage dense tensor cores for small block sizes. This design choice supports fine-grained KV-cache pruning.
Table 9: FlashInfer Fine-Grained Sparsity Latency (us)
page_budget
seq_len 64 128 256 512
4096 20.299 30.361 44.383 44.430
8192 22.273 28.603 44.928 68.194
16384 20.485 28.678 44.677 68.700
32768 22.371 28.700 44.988 68.478
Table 10: PyTorch SDPA Fine-Grained Sparsity Latency (us)
page_budget
seq_len 64 128 256 512
4096 287.684 288.904 287.715 287.807
8192 474.631 474.508 474.683 473.070
16384 857.319 857.570 857.094 857.728
32768 1711.955 1711.621 1713.093 1711.709
Table 11: FlexAttention Fine-Grained Sparsity Latency (us)
page_budget
seq_len 64 128 256 512
4096 1100.349 1097.356 1073.753 1071.797
8192 1092.695 1099.100 1078.081 1074.886
16384 1109.817 1101.535 1077.639 1076.859
32768 1169.109 1187.395 1176.332 1174.502